Consider the following equilibrium.

A+2B⇌C+3D

If the change in concentration for C is 0.760 M, what will be the corresponding change in concentration (in units of molarity) for D?

  • Round the answer to three significant figures.
Answer #1

Calculation:
Given the equilibrium reaction:

A+2BC+3D

The stoichiometry shows that for every 1 mole of C produced, 3 moles of D are produced.


  • Change in concentration for C (Δ[C]) = +0.760 M (since it is formed).


  • Change in concentration for D (Δ[D]) = 3×Δ[C]=3×0.760M=2.28M.


Thus, the change in concentration for D is 2.28 M.


Answer:
The corresponding change in concentration for D is 2.28 M.


answered by: Harshwardhan kunal